Draw one card from a well - shuffled deck of 52 playing cards. What is the probability of drawing a diamond card

Answer by mananth(16949)

52 cards in a pack of cards
diamonds there are 13 in a pack
p(D) = 13/52 = 1/4
probability of drawing a diamond card
=1/4